# A theoretical analysis of mass scaling techniques

### v1

@@ 3.2 Local deflation
Choosing a large value of $\alpha$ may quickly undermine the quality of the solution as it moves inaccurate high frequency modes to the low frequency regime. González and Park [23] already highlighted the limitations of this technique for a simple bar model and its alarming impact on the accuracy. Thus, one should exercise caution when using this technique. The second strategy, on the contrary, always preserves the eigenvalue ordering. The result, illustrated in fig. 3.9 for $r=12$, consists in shaving off the upper part of the spectrum.
@@
We now investigate the global properties of the method using the same finite element mesh and material properties as in the previous section. Figure 10(a) shows the increase of the critical time step for the first strategy for fixed rank values and increasing values of $\alpha$ together with the upper bound from corollary 2.17. Interestingly, fig. 10(a) displays a threshold effect reminiscent of the local behavior. Figure 10(b) shows the increase of the critical time step for the second strategy, which is perfectly captured by the upper bound of corollary 2.18.
@@
The eigenfrequency ratio for the entire spectrum is shown in fig. 3.11 for both strategies and selected values of scaling parameter $\alpha$ and rank $r$. Contrary to the method of Olovsson et al., the lower bound is sharp only for the largest eigenfrequencies, which is the desired behavior.
@@ 3.3 Summary
The ad hoc mass scaling strategies of Olovsson et al. and Hoffmann et al. may be viewed as inexact local deflation strategies based on low-frequency modes. One notable advantage of using low-frequency modes instead of high-frequency ones is that the former includes rigid-body modes and are known regardless of how distorted the element is. Thus, they offer better guarantees of robustness. Moreover, engineers have developed a very good understanding of the deformation mode shapes for an hexahedral element (see [25, Figure 3]), based on which they constructed these ad hoc strategies. While the method of Hoffmann et al. seems significantly better than the one of Olovsson et al., it still does not exactly target the right modes, which both impedes on the accuracy and the increase of the critical time step. Preserving the eigenvalue ordering seems essential to avoid interchanging modes, which might disastrously affect the accuracy of the solution. Exact local deflation strategies may deliver far better results but assume that the stiffness matrix is explicitly available, which is rarely the case in commercial software, especially for nonlinear problems. We believe it might be possible to further improve the method of Hoffmann et al. for obtaining the desired behavior but leave this task to the engineering community.

### v2

@@ 3.2 Local deflation
Choosing a large value of $\alpha$ may quickly undermine the quality of the solution as it moves inaccurate high frequency modes to the low frequency regime. González and Park [25] already highlighted the limitations of this technique for a simple bar model and its alarming impact on the accuracy. Thus, one should exercise caution when using this technique. The second strategy, on the contrary, always preserves the eigenvalue ordering. The result, illustrated in Figure 3.9 for $r=12$, consists in shaving off the upper part of the spectrum.
@@
We now investigate the global properties of the method using the same finite element mesh and material properties as in the previous section. Figure 10(a) shows the increase of the critical time step for the first strategy for fixed rank values and increasing values of $\alpha$ together with the upper bound from Corollary 2.17. Interestingly, Figure 10(a) displays a threshold effect reminiscent of the local behavior. Figure 10(b) shows the increase of the critical time step for the second strategy, which is perfectly captured by the upper bound of Corollary 2.18.
@@
The eigenfrequency ratio for the entire spectrum is shown in Figure 3.11 for both strategies and selected values of scaling parameter $\alpha$ and rank $r$. Contrary to the method of Olovsson et al., the lower bound is sharp only for the largest eigenfrequencies, which is the desired behavior.
@@ 3.3 Summary
The ad hoc mass scaling strategies of Olovsson et al. and Hoffmann et al. may be viewed as inexact local deflation strategies based on low-frequency modes. One notable advantage of using low-frequency modes instead of high-frequency ones is that the former includes rigid-body modes and are known regardless of how distorted the element is. Thus, they offer better guarantees of robustness. Moreover, engineers have developed a very good understanding of the deformation mode shapes for an hexahedral element (see [27, Figure 3]), based on which they constructed these ad hoc strategies. While the method of Hoffmann et al. seems significantly better than the one of Olovsson et al., it still does not exactly target the right modes, which both impedes on the accuracy and the increase of the critical time step. Preserving the eigenvalue ordering seems essential to avoid interchanging modes, which might disastrously affect the accuracy of the solution. Exact local deflation strategies may deliver far better results but assume that the stiffness matrix is explicitly available, which is rarely the case in commercial software, especially for nonlinear problems. We believe it might be possible to further improve the method of Hoffmann et al. for obtaining the desired behavior but leave this task to the engineering community.
@@ 4 Conclusion
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.